| Name | Email | Date | Comment
|
| phil
| phil@naic.edu
| 06/12/2009
| installed new feed, filters before dewar. Dewar is still room temperature.
|
| phil
| phil@naic.edu
| 06/15/2009
| replaced cables feed to dewar with good cables. Installed filters after the dewar. Retuned dipoles.
|
| phil
| phil@naic.edu
| 06/29/2009
| installed cooled receiver.
|
| dana | dwhitlow@naic.edu | 07/03/2009 | late 02july or early 03jul09:
New (lower loss) CAL couplers were installed between feed and mini-dewar input. I believe the couplers were installed between the large filters and the dewar.
|
| crv | | Thursday, August 20, 2009 15:15:45 | fuse was blowned and the reciever warmed on the night of tuesday the fuse was replaced and the reciever cooldown was restarted.Reciever is cold.
|
| crv | | Tuesday, December 22, 2009 15:17:12 | Penn state feed Antenna has been replaced by a BYU feed antenna that will be tested for a couple of days. This antenna has only two of the four dipole connected(3,4). These dipoles are connected to Ch A.
|
| crv | crios@nai.edu | Wednesday, February 24, 2010 16:56:13 | The system is warm since Feb.10th 2010.
|
| crios | | Friday, April 30, 2010 17:31:55 | On April 23rd the power cord of cryo pump was replaced and started the cooldown.
|
| crv | crios@naic.edu | Wednesday, June 2, 2010 10:44:40 | The system was turned off on May 27th,2010. The system will be lowered on the week to make space for the BYU proyect.
|
| crv | crios@naic.edu | Monday, September 27, 2010 13:36:43 | The system was installed and started cooling on friday sep. 27 2010 and is ready for testing.
|
| crv | | Monday, September 27, 2010 17:01:20 | System feed array is BYU 1st desing.
|
| crv | crios@naic.edu | Thursday, October 28, 2010 13:14:23 | The reciever is not available for use since yesterday due to problem with the dewar cryo pump and power supply.
|
| crv | crios@naic.edu | Friday, November 5, 2010 10:56:53 | The dewar was lowered to the shop to be repaired.
|
| crv | crios@naic.edu | Wednesday, November 17, 2010 14:37:35 | The system has been repaired and is under test by Dana and is schedule to be installed tomorrow.
|
| crv | crios@naic.edu | Thursday, November 18, 2010 14:04:36 | The system was installed today and started cooldown and its 77.8K now.
|
| GR | grivera@naic.edu | Monday, November 29, 2010 13:08:32 | The coldhead fuse was blown yesterday (Sunday, Nov. 28). The fuse was replaced and it was left cooling.
|
| crv | crios@naic.edu | Friday, January 14, 2011 16:21:13 | System started cooldown at 10:00am.
|
| crv | | Monday, February 14, 2011 14:16:38 | The reciever was turned off .because it warmed up do to a fualt on the vac. pump. The dewar has a new vac. pump connected and will be pumpping over night.
|
| crios | | Wednesday, May 25, 2011 15:33:15 | The cryo temp. was 140K Denis and I found that the vac. pump was off this happened after several power dips.
The dewar is cooling down now and the vac. was .6microns.
|
| Denis Urbain | durbain@naic.edu | Wednesday, April 3, 2013 13:26:20 | The receiver was dismantled and will be removed from the rotary floor. We need the space for the upcoming AO19 Rx test, plus the Rx has not been requested for a long time.
|